OpenStudy (anonymous):

What is the sum of 3/2x and 4/3x expressed in simplest form?

OpenStudy (angela210793):

3/2x+4/3x=(9x+8x)/6x^2=17x/6x^2

OpenStudy (anonymous):

how do you get the 17?

OpenStudy (anonymous):

(3x/2)+(4x/3) (9x+8x)6 17x/6

OpenStudy (angela210793):

it's 6x^2 Math

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

How did you guys get the 17?

OpenStudy (anonymous):

(3/2)x+(4/3)x

OpenStudy (angela210793):

Ajmal....don't u have 2x and 3x in the denoms? so 2x*3x=6x^2

OpenStudy (anonymous):

literally none of these answers are helping me as to how you got 17.. i'm lost.

OpenStudy (angela210793):

17=8+9

OpenStudy (anonymous):

how'd you get 8 and 9?

OpenStudy (anonymous):

OH NEVERMIND! I GOT IT NOW! hahaha i feel like an idiot! thanks for your help guys! :)

OpenStudy (angela210793):

uw :D

OpenStudy (anonymous):

rebecca your question is not clear (3/2)x+(4/3)x or 3/2x+4/3x

OpenStudy (angela210793):

yea actually Math. is right....